# ConsDreamer: Advancing Multi-View Consistency for Zero-Shot Text-to-3D Generation

### v1

@@ Abstract
Recent advances in zero-shot text-to-3D generation have revolutionized 3D content creation by enabling direct synthesis from textual descriptions. While state-of-the-art methods leverage 3D Gaussian Splatting with score distillation to enhance multi-view rendering through pre-trained text-to-image (T2I) models, they suffer from inherent view biases in T2I priors. These biases lead to inconsistent 3D generation, particularly manifesting as the multi-face Janus problem, where objects exhibit conflicting features across views. To address this fundamental challenge, we propose ConsDreamer, a novel framework that mitigates view bias by refining both the conditional and unconditional terms in the score distillation process: (1) a View Disentanglement Module (VDM) that eliminates viewpoint biases in conditional prompts by decoupling irrelevant view components and injecting precise camera parameters; and (2) a similarity-based partial order loss that enforces geometric consistency in the unconditional term by aligning cosine similarities with azimuth relationships. Extensive experiments demonstrate that ConsDreamer effectively mitigates the multi-face Janus problem in text-to-3D generation, outperforming existing methods in both visual quality and consistency.
Index Terms:
text to 3D generation, multi-face Janus problem, score distillation sampling.
@@ I Introduction
[lines=2]3D GENERATION technology plays a crucial role in various fields such as innovative industrial design, game development, and virtual reality. In particular, zero-shot text-to-3D generation [1], [2], [3], [4], [5] aims to generate 3D content without 3D training data, enabling the conversion from concept to reality. However, zero-shot text-to-3D generation tasks [6], [7], [8], [9] are constrained by the inherent complexity of the wild world and the scarcity of 3D data, unlike text-to-image (T2I) tasks [10], [11]. From this perspective, generating high-quality 3D content from text is still a significant challenge.
Earlier works [12], [13], [14], [15], [16] based on Neural Radiance Fields (NeRF) introduce a framework that extended the 2D generative capabilities of diffusion models into the 3D domain by leveraging a view-conditioned projection mechanism. For instance, DreamFusion [14] introduces the Score Distillation Sampling (SDS) technique, effectively aligning the rendering of 3D content with 2D generative priors. However, rendering in NeRF primarily relies on implicit 3D representations, which significantly increases computational cost and makes it challenging to produce high-resolution images [17]. Recently, 3D Gaussian Splatting [18] has been introduced as an efficient alternative to NeRF’s implicit 3D representation, enabling real-time, high-resolution rendering through explicit point-based modelling. Its adoption in SOTA zero-shot text-to-3D generation methods [17], [19], [20], [21] highlights its effectiveness in delivering superior results in quality and speed. However, these recent methods still struggle with generation precision, e.g., the multi-face Janus problem [13], [14]. A single 3D object exhibiting inconsistencies across its faces in different views, as shown in Fig. 1, significantly undermines the realism and coherence of the 3D output. Therefore, ensuring view consistency in 3D generation is crucial for high-quality results, which is the primary motivation for this paper.
@@
Figure 1: Examples of Multi-Face Janus Problem: (a) generated by DreamFusion [14], (b) by Sherpa3D [22], (c) by LucidDreamer [17], (d) by DreamScene [23], and (e) by ScaleDreamer [24].
@@
Figure 2: Examples of text-to-3D content creation using our framework. We present a text-to-3D generation framework, named $ConsDreamer$, which leverages a View Disentanglement Model and a novel partial order loss to ensure semantic clarity across views (detailed in Section III). The generative 3D results demonstrate the superiority of ConsDreamer. Please zoom in for details.
@@
Considering the most recent text-to-3D methods, they predominantly rely on T2I models to generate reference images. However, these models are typically trained on datasets sourced from publicly available resources and online platforms, such as photo-sharing sites. A notable limitation of these datasets is the lack of diverse and well-annotated multi-view images, which results in a bias toward high-frequency single view text-image pairs. As a consequence, T2I models tend to prioritise certain viewpoints, particularly frontal or side views [25]. This view bias adversely impacts the consistency of 3D generation. When given rendered images and corresponding textual descriptions specifying particular views, the model often defaults to its prior view preference, rather than adhering to the exact view that training required. This misalignment leads to multi-view inconsistencies, where features across different views are incompatible, exacerbating the multi-face Janus problem in text-to-3D generation. Furthermore, the dominance of preferred viewpoints introduces an unbalanced view distribution, where even in the absence of explicit view conditions, the model still favours certain perspectives.
To tackle the aforementioned problem, we first conducted a mathematical analysis of the Janus problem. The analysis reveals that view biases arise during parameter updates, affecting both conditional and unconditional terms used in existing text-to-3D methods. In conditional terms, the input prompts guide view control, but conflicts between prior view preferences and target view specifications result in ambiguous view semantics. Unconditional terms, independent of prompts, are directly shaped by inherent view biases. Therefore, this paper proposes a novel framework, ConsDreamer, to address the challenges from both conditional and unconditional perspectives. For the conditional term, we propose the View Disentanglement Module (VDM), which starts extracting view-agnostic keyword embeddings from the prompt by combining the extracted keyword with explicit view descriptions (e.g., “front view,” “side view”). These view-specific embeddings are processed to isolate view-related features by subtracting their projection from the view-agnostic embedding space. The resulting view-specific features are used to eliminate prior view preferences and inject target view information, ensuring that the generated content adheres to user-specified views.
@@
For the unconditional term, we introduce a similarity-based partial order loss $\mathcal{L}_{P}$, which leverages the inherent relationship between view similarity and azimuthal angle distances. To implement this, a Cartesian coordinate framework is established, and an expected similarity partial order is determined based on the proximity of the azimuthal angles of multiple views to a reference view. The actual cosine similarity scores of the rendered images are constrained to conform to this partial order. By enforcing these constraints, $\mathcal{L}_{P}$ ensures cross-view consistency in 3D content, effectively addressing the multi-face Janus problem. We also conducted comprehensive experiments to illustrate the efficacy of our method, confirming that ConsDreamer can better alleviate the multi-face Janus problem and strengthen semantic correlations across different views and enhance generated content quality. The main contributions of this work are summarized as follows:
@@
    We conduct extensive experiments on a wide range of prompts to validate the effectiveness of the proposed method.
@@ II-A Differentiable 3D Representations
Differentiable 3D representation methods are especially essential for zero-shot text-guided 3D generation. Their differentiable nature enables a trainable pathway, optimizing 3D parameters by ensuring that rendered results from random views align with the prompt inputs. Previously, various representations are introduced for text-to-3D generation [12], [26], [27], [28]. Among these, NeRF [12] emerges as the most widely adopted representation in text-to-3D generation tasks. NeRF enables novel view synthesis from unobserved viewpoints and leverages guidance information from the T2I model to optimize the tri-plane [29] parameters through a pre-trained MLP network, demonstrating broad applicability across downstream 3D tasks [3], [12], [13], [14], [15]. Despite its success, NeRF faces challenges in rendering speed and memory consumption due to its implicit nature. A new 3D representation method, named 3D Gaussian Splatting, is thus introduced. Gaussian Splatting-based methods [18], [17], [30], [22] address these limitations using explicit 3D representations with anisotropic Gaussians. They ensure high reconstruction quality, real-time rendering, and reduced computational demands, making them ideal for large-scale and dynamic scenes. By replacing the NeRF method with the explicit 3D Gaussian Splatting representation, it becomes possible to integrate pre-trained 3D point cloud models like Point-E [31] and Shape-E [32] into the 3D generation framework, providing a prior 3D knowledge foundation.
II-B Diffusion models
Diffusion models serve as a core component for guiding text-to-3D generation. The Denoising Diffusion Probabilistic Model (DDPM) [33], based on Markov chains, operates by gradually “diffusing” data from its original state into random noise in the forward procedure and then reversing the process to reconstruct the original data. Although DDPM achieves “diffusion-reconstruction,” its efficiency is hindered by the step-by-step sampling strategy. Subsequently, Denoising Diffusion Implicit Models (DDIM) [11] introduce a deterministic sampling approach, enabling high-quality generation with significantly fewer sampling steps. Additionally, DDPM faces high time consumption as the data retains its original size at every step. Latent Diffusion Models (LDM) [34] address this limitation by transferring the diffusion process from the original pixel space to a low-dimensional latent space, significantly improving efficiency, where the attention [7], [35] mechanism could help to control the learning process. Most recently, Stable Diffusion [34] introduces a conditional generation approach based on the LDM, enabling high-quality image generation from textual descriptions. Subsequent research [17], [36], [37] extends Stable Diffusion with the deterministic sampling mechanism in DDIM.
II-C Text-to-3D Generation
@@
Early efforts, such as DreamField [38], leverage the cross-modal information of CLIP [39] to translate text into 3D. However, DreamField faces limitations due to the lack of high-precision alignment in CLIP’s cross-modal information, resulting in suboptimal 3D generation performance. With the advent of diffusion models, a new paradigm [14], [30] emerges for extracting 3D assets from pre-trained T2I models, leveraging 2D diffusion model outputs to guide differentiable 3D representations. For example, DreamFusion introduces Score Distillation Sampling (SDS), which aligns 3D rendering information with diffusion priors and optimizes model parameters through backpropagation. However, SDS-based approaches often introduce smoothing effects. Recent research [40], [41], [42], [43] aims to address this. Prolific-Dreamer [41] proposes Variational Score Distillation (VSD), significantly improving generation quality. Consistent 3D [9] analyzes SDS from ordinary differential equations (ODEs) and introduces Consistency Distillation Sampling (CSD). ScaleDreamer [24] introduces Asynchronous Score Distillation (ASD), which shifts diffusion timesteps to minimize noise prediction errors without fine-tuning the diffusion model. LucidDreamer [17] addresses over-smoothing issues via Interval Score Matching (ISM), integrating DDIM inversion to enable deterministic sampling. Subsequent advancements, DreamerXL [37], ExactDreamer [36], and Guided Consistency Sampling (GCS) [21] further refine this paradigm by systematically reducing accumulated errors in DDIM’s reverse process. The methods above are predominantly optimization-based, while recent advances also explore feed-forward methodologies [44], [45], [46] to facilitate the generation of large-scale 3D assets efficiently [47] or to integrate auxiliary information such as text-to-3D knowledge graphs [48]. These methods, however, demand substantial GPU resources for additional training and require extensive 3D datasets.
@@
Figure 3: An overview of ConsDreamer. Our framework is built upon the main flow of 3D content distilled from the T2I model. ConsDreamer introduces two key innovations: (a) VDM disentangles the keyword in the prompt to obtain the canonical view features $\delta_{\text{key}}^{\text{view}}$, which are then used for precise view control (detailed in Section III-B). (b) A novel partial order loss $\mathcal{L}_{P}$ is introduced among multi-view rendered images to endow the model with view-aware capabilities (detailed in Section III-C). Together, the VDM and $\mathcal{L}_{P}$ enhance the clarity of view semantics and significantly mitigate the multi-face Janus problem.
Several works specifically address the multi-face Janus problem. Perp-Neg [25] leverages the geometric properties of score space to enhance the utilization of negative prompts. MVDream [13] fine-tunes pre-trained T2I models using canonical view image combinations, ensuring view consistency but requiring additional training and datasets. Debiased-SDS [15] mitigates the multi-face Janus problem in two ways: first, by clipping distillation scores to suppress erroneous artifacts, and second, by identifying and omitting words in the prompt that conflict with the target view. However, Debiased-SDS lacks fine-grained view control, and omitting conflicting words can compromise the integrity of user prompts. Sherpa3D [22] uses structural alignment and semantic guidance, incorporating normal map gradients and CLIP semantic embeddings to preserve the geometric consistency of the initial 3D prior during 2D lifting optimization. However, this method heavily relies on coarse 3D prior information. Overall, existing approaches lack precise understanding and control of view-related information while requiring additional computational resources. In this paper, we introduce a novel framework that eliminates prior view biases by disentangling view features and ensuring multi-view consistency through a carefully designed partial order loss, significantly enhancing view semantic clarity without requiring additional training.

### v2

@@ Abstract
Recent advances in zero-shot text-to-3D generation have revolutionised 3D content creation by enabling direct synthesis from textual descriptions. While state-of-the-art methods leverage 3D Gaussian Splatting with score distillation to enhance multi-view rendering through pre-trained text-to-image (T2I) models, they suffer from inherent prior view biases in T2I Models. These biases lead to inconsistent 3D generation, particularly manifesting as the multi-face Janus problem, where objects exhibit conflicting features across views. To address this fundamental challenge, we propose ConsDreamer, a novel method that mitigates view bias by refining both the conditional and unconditional terms in the score distillation process: (1) a View Disentanglement Module (VDM) that eliminates viewpoint biases in conditional prompts by decoupling irrelevant view components and injecting precise view control; and (2) a similarity-based partial order loss that enforces geometric consistency in the unconditional term by aligning cosine similarities with azimuth relationships. Extensive experiments demonstrate that ConsDreamer can be seamlessly integrated into various 3D representations and score distillation paradigms, effectively mitigating the multi-face Janus problem.
@@ I Introduction
3DGENERATION technology plays a crucial role in various fields such as innovative industrial design, game development, and virtual reality. In particular, zero-shot text-to-3D generation [chen2023fantasia3d-lucid-5], [hong2022avatarclip-lucid-14], [lin2023magic3d-lucid-21], [metzer2023latent-lucid-29], [michel2022text2mesh-lucid-31] aims to generate 3D content without 3D training data, enabling the conversion from concept to reality. However, zero-shot text-to-3D generation tasks [chen2019text2shape-2405-1], [duan2023dynamic-2405-4], [duan2020sofa-2405-5], [wu2016learning-2405-40] are constrained by the inherent complexity of the wild world and the scarcity of 3D data, unlike text-to-image (T2I) tasks [ruiz2023dreambooth-2405-30], [saharia2022photorealistic-2405-32]. From this perspective, generating high-quality 3D content from text is still a significant challenge.
@@
Figure 1: Examples of the Multi-Face Janus Problem: (a) generated by DreamFusion [poole2022dreamfusion-lucid-34], (b) by SDI [lukoianov2024score], (c) by LucidDreamer [liang2024luciddreamer-lucid], (d) by DreamScene [10.1007/978-3-031-72904-1_13-dreamscene], and (e) by GE3D [li2024text].
Earlier works [mildenhall2021nerf-lucid-32], [shi2023mvdream-lucid-41], [poole2022dreamfusion-lucid-34], [hong2023debiasing-disbais], [miao2025laser] based on Neural Radiance Fields (NeRF) introduce a framework that extended the 2D generative capabilities of diffusion models into the 3D domain by leveraging a view-conditioned projection mechanism. For instance, DreamFusion [poole2022dreamfusion-lucid-34] introduces the Score Distillation Sampling (SDS) technique, effectively aligning the rendering of 3D content with 2D generative priors. However, rendering in NeRF primarily relies on implicit 3D representations, which significantly increases computational cost and makes it challenging to produce high-resolution images [liang2024luciddreamer-lucid]. Recently, 3D Gaussian Splatting [kerbl20233d-2405-13] has been introduced as an efficient alternative to NeRF’s implicit 3D representation, enabling real-time, high-resolution rendering through explicit point-based modelling. Its adoption in State-of-the-Art (SOTA) zero-shot text-to-3D generation methods [liang2024luciddreamer-lucid], [sun2024gseditpro-3dgswork1], [di2024hyper-hyper3d], [10.1007/978-3-031-72775-7_16-gcs] highlights its effectiveness in delivering superior results in quality and speed. However, these recent methods still struggle with generation precision, e.g., the multi-face Janus problem [shi2023mvdream-lucid-41], [poole2022dreamfusion-lucid-34]. A single 3D object exhibiting inconsistencies across its faces in different views, as shown in Fig. 1, significantly undermines the realism and coherence of the 3D output. Therefore, ensuring view consistency in 3D generation is crucial for high-quality results, which is the primary motivation for this paper.
@@
Figure 2: Examples of text-to-3D content creation using our method. We introduce ConsDreamerConsDreamer, a method that guarantees multi-view consistency by leveraging a View Disentanglement Model and a novel partial order loss to ensure semantic clarity across views
(detailed in Section III). The generative 3D results demonstrate the superiority of ConsDreamer. Please zoom in for details.
@@
Considering the most recent text-to-3D methods, they predominantly rely on T2I models to generate reference images. However, these models are typically trained on datasets sourced from publicly available resources and online platforms, such as photo-sharing sites. A notable limitation of these datasets is the lack of diverse and well-annotated multi-view images, which results in a bias toward high-frequency single view text-image pairs. As a consequence, T2I models tend to prioritise certain viewpoints, particularly frontal or side views [armandpour2023re-perpneg]. This view bias adversely impacts the consistency of 3D generation. When given rendered images and corresponding textual descriptions specifying particular views, the model often defaults to its prior view preference, rather than adhering to the exact view that training required. This misalignment leads to multi-view inconsistencies, where features across different views are incompatible, exacerbating the multi-face Janus problem in text-to-3D generation. Furthermore, the dominance of preferred viewpoints introduces an unbalanced view distribution, where even in the absence of explicit view conditions, the model still favours certain perspectives.
Figure 3: Comparison of PerpNeg and VDM in handling prior view biases. (a) Perp-Neg works after UNet denoising in the score space by orthogonalising negative-view scores. However, frontal-view features that have already been entangled into the subject-keyword CA map are absorbed into the preserved component $\textit{neg}_{\parallel}$, so prior-view faces survive and still cause Janus artefacts. (b) Our VDM instead intervenes before denoising in the prompt-embedding space: it extracts view-specific residuals from view-augmented prompts, subtracts prior-view residuals, and injects only the residual corresponding to the target azimuth, so the conditioning entering the UNet is already view-corrected, producing clean CA maps and substantially reducing Janus.
To tackle the aforementioned problem, we first conduct a mathematical analysis of the Janus issue. This analysis reveals that the prior view biases in T2I models affect both the conditional and unconditional terms used in existing text-to-3D methods. The unconditional term, which is independent of the textual prompt, is directly shaped by these inherent view biases. In the conditional term, prompt embeddings are fused with the target view description and then fed into the UNet as conditioning signals to generate the specified view. However, when the prior view preference of the T2I model disagrees with the desired target view, their interaction leads to ambiguous view semantics and unstable gradients. To alleviate this problem, Perp-Neg [armandpour2023re-perpneg] introduces a prompt-handling strategy that orthogonalises view-irrelevant components in the denoising score space, and has been widely adopted in recent text-to-3D pipelines. In Perp-Neg, the score corresponding to the target view is treated as the positive condition pos, while the score associated with an irrelevant view is treated as the negative condition neg. The method computes an update direction from neg to pos, and decomposes neg into a component parallel to pos, $\textit{neg}_{\parallel}$, and an orthogonal component, $\textit{neg}_{\perp}$. The parallel part $\textit{neg}_{\parallel}$ is retained to preserve subject semantics, while $\textit{neg}_{\perp}$ is interpreted as view-specific content and is suppressed. However, when the target view conflicts with prior view preference, the cross-attention (CA) map of the subject-keyword (e.g., “squirrel”) still exhibits strong frontal-view activations. These erroneous activations are then incorrectly absorbed into $\textit{neg}_{\parallel}$ as subject semantics and thus preserved, which allows the multi-face Janus artefacts to persist. The distortions observed in the conditional results closely follow the spatial distribution of these abnormal facial activations, as illustrated in Fig. 3 (a). Therefore, we propose a View Disentanglement Module (VDM) that removes prior view bias and explicitly injects target-view control before the denoising process, thereby avoiding the accumulation of prior-view features inside the UNet. VDM computes subject-keyword view-specific residuals from the original prompt and the view-augmented prompt, and the residuals are then used to cancel prior view bias and to inject the desired target-view information according to the camera azimuth. In this way, the conditioning representation entering the UNet already aligns with the user-specified view, leading to cleaner CA maps and substantially reducing Janus artefacts in multi-view text-to-3D generation.
@@
For the unconditional term, we introduce a similarity-based partial order loss $\mathcal{L}_{\text{P}}$, which leverages the inherent relationship between view similarity and azimuthal angle distances. To implement this, a Cartesian coordinate framework is established, and an expected similarity partial order is determined based on the proximity of the azimuthal angles of multiple views to a reference view. The actual cosine similarity scores of the rendered-image features are constrained to conform to this partial order. By enforcing these constraints, $\mathcal{L}_{\text{P}}$ ensures cross-view consistency in 3D content, effectively addressing the multi-face Janus problem. We also conducted comprehensive experiments to illustrate the efficacy of our method, ConsDreamer, which confirms that it can better alleviate the multi-face Janus problem, strengthen inter-view semantic correlations, and enhance the quality of generated content. The main contributions of this work are summarized as follows:
@@
    We conduct extensive experiments on a wide range of prompts to verify the effectiveness of the proposed method when integrated into different 3D representations and various score-distillation strategies.
@@ II-A Differentiable 3D Representations
Differentiable 3D representation methods are especially essential for zero-shot text-guided 3D generation. Their differentiable nature enables a trainable pathway, optimizing 3D parameters by ensuring that rendered results from random views align with the prompt inputs. Previously, various representations are introduced for text-to-3D generation [mildenhall2021nerf-lucid-32], [barron2021mip-lucid-3], [ge2023ref-lucid-8], [shen2021deep-lucid-40]. Among these, NeRF [mildenhall2021nerf-lucid-32] emerges as the most widely adopted representation in text-to-3D generation tasks. NeRF enables novel view synthesis from unobserved viewpoints and leverages guidance information from the T2I model to optimize the tri-plane [shue20233d-triplane] parameters through a pre-trained MLP network, demonstrating broad applicability across downstream 3D tasks [lin2023magic3d-lucid-21], [mildenhall2021nerf-lucid-32], [shi2023mvdream-lucid-41], [poole2022dreamfusion-lucid-34], [hong2023debiasing-disbais]. Despite its success, NeRF faces challenges in rendering speed and memory consumption due to its implicit nature. A new 3D representation method, named 3D Gaussian Splatting, is thus introduced. Gaussian Splatting-based methods [kerbl20233d-2405-13], [liang2024luciddreamer-lucid], [tang2023dreamgaussian-lucid-45], [liu2024sherpa3d-sherpa3d] address these limitations using explicit 3D representations with anisotropic Gaussians. They ensure high reconstruction quality, real-time rendering, and reduced computational demands, making them ideal for large-scale and dynamic scenes. By replacing the NeRF method with the explicit 3D Gaussian Splatting representation, it becomes possible to integrate pre-trained 3D point cloud models like Point-E [nichol2022point-lucid-33] and Shape-E [jun2023shap-sherpa3d-28] into the 3D generation framework, providing a prior 3D knowledge foundation.
II-B Text-to-3D Generation
@@
Early efforts, such as DreamField [jain2022zero-2405-10], leverage the cross-modal information of CLIP [ramesh2021zero-2405-26] to translate text into 3D. However, DreamField faces limitations due to the lack of high-precision alignment in CLIP’s cross-modal information, resulting in suboptimal 3D generation performance. With the advent of diffusion models, a new paradigm [poole2022dreamfusion-lucid-34], [tang2023dreamgaussian-lucid-45] emerges for extracting 3D assets from pre-trained T2I models, leveraging 2D diffusion model outputs to guide differentiable 3D representations. For example, DreamFusion introduces Score Distillation Sampling (SDS), which aligns 3D rendering information with diffusion priors and optimizes model parameters through backpropagation. However, SDS-based approaches often introduce smoothing effects. Recent research [katzir2023noise-lucid-18], [wang2024prolificdreamer-lucid-47], [yu2023text-lucid-49], [zhu2023hifa-lucid-52] aims to address this. Prolific-Dreamer [wang2024prolificdreamer-lucid-47] proposes Variational Score Distillation (VSD), significantly improving generation quality. Consistent 3D [wu2016learning-2405-40] analyzes SDS from ordinary differential equations (ODEs) and introduces Consistency Distillation Sampling (CSD). ScaleDreamer [10.1007/978-3-031-72667-5_1-scaledreamer] introduces Asynchronous Score Distillation (ASD), which shifts diffusion timesteps to minimize noise prediction errors without fine-tuning the diffusion model. LucidDreamer [liang2024luciddreamer-lucid] addresses over-smoothing issues via Interval Score Matching (ISM), integrating Denoising Diffusion Implicit Models (DDIM) [saharia2022photorealistic-2405-32] inversion to enable deterministic sampling. Subsequent advancements, DreamerXL [miao2024dreamer-dreamxl], ExactDreamer [zhang2024exactdreamer-exactdreamer], and Guided Consistency Sampling (GCS) [10.1007/978-3-031-72775-7_16-gcs] further refine this paradigm by systematically reducing accumulated errors in DDIM’s reverse process. The methods above are predominantly optimization-based, while recent advances also explore feed-forward methodologies [xu2023dmv3d], [liu2024pi3d], [cao2023large-difftf] to facilitate the generation of large-scale 3D assets efficiently [cao2025difftf++] or to integrate auxiliary information such as text-to-3D knowledge graphs [nie2024t2td-t2td]. These methods, however, demand substantial GPU resources for additional training and require extensive 3D datasets.
II-C Mitigating the Multi-Face Janus Issue
Several studies have specifically tackled the multi-face Janus problem. Among these, Perp-Neg [armandpour2023re-perpneg] stands out as the most widely adopted, which leverages the geometric properties of the score space to enhance the utilisation of negative prompts. However, this method operates solely on the UNet output, thereby neglecting the highly aggregated prior view information throughout the UNet processing pipeline. Consequently, the multi-face Janus problem may still recur when lifting 2D representations to 3D. MVDream [shi2023mvdream-lucid-41], Zero123 [liu2023zero1to3], and SV3D [voleti2024sv3d] all rely on pre-trained models and finetuning, which introduces high computational costs and dependency on large datasets. While MVDream ensures view consistency through fine-tuning, and Zero123 and SV3D generate multi-view 3D representations, these methods often struggle with maintaining consistent 3D geometry and texture across views. Moreover, they require extensive training, making them less efficient for quick deployment. In this paper, we introduce a novel method that eliminates prior view biases by disentangling view features and ensuring multi-view consistency through a carefully designed partial order loss. This approach significantly enhances view semantic clarity without the need for additional training or 3D datasets, offering a versatile, rapid-deployment solution that can be easily applied to a wide range of 3D generation frameworks.
@@ II-C Mitigating the Multi-Face Janus Issue
Figure 4: An overview of ConsDreamer. Our method is built upon the main flow of 3D content distilled from the T2I model. ConsDreamer introduces two key innovations: (a) VDM disentangles the keyword in the prompt to obtain the canonical view features $\delta_{\text{key}}^{\text{view}}$, which are then used for precise view control (detailed in Section III-B). (b) A novel partial order loss $\mathcal{L}_{\text{P}}$ is introduced among multi-view rendered images to endow the model with view-aware capabilities (detailed in Section III-C). Together, the VDM and $\mathcal{L}_{\text{P}}$ enhance the clarity of view semantics and significantly mitigate the multi-face Janus problem.